OpenStudy (anonymous):

What is the sum of a 32–term arithmetic sequence where the first term is 12 and the last term is 105

OpenStudy (anonymous):

sum = n*(a1+an)/2 = 32*(12+105)/2 =....

OpenStudy (anonymous):

1872 i guess

OpenStudy (anonymous):

welll your guesses are pretty good so i will take it hhaha thanks yo!!(:

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Can you help with this one? What is the sum of a 17–term arithmetic sequence where the first term is 9 and the last term is –87?

OpenStudy (anonymous):

what??? it's the same it's 17*(9-87)/2 = -663 i guess lol

OpenStudy (anonymous):

whyyy are you sooo smart! haha it is, what about What is the sum of the arithmetic sequence 6, 11, 16 …, if there are 24 terms?

OpenStudy (anonymous):

haha